OpenStudy (anonymous):

Solve the system and explain what the results mean according to your scenario. 2x + y = 9 3x + 4y = 26

OpenStudy (lilg132):

2x + y = 9 3x + 4y = 26 2x + y = 9 (multiply by 3) 3x + 4y = 26 (multiply by 2) so we have the same value for x and we can cancel it out 6x + 3y = 27 6x + 8y = 52 5y = 25 y = 5 now use this to work out x

OpenStudy (lilg132):

so just so you understand i will show you how to 2x + y = 9 3x + 4y = 26 so we know the value of y = 5 2x + 5 = 9 rearrange 2x = 9 - 5 2x = 4 x = 2 to check use these answers in the second equation x = 2 y = 5 3(2) + 4(5) = 26 6 + 20 = 26 dont forget to click good answer :)

OpenStudy (anonymous):

you can use elimination of substitution. in this case it's easier to use substitution. subtract 2x from both sides of the first equation to get the y alone y = 9 - 2x now, you want to plug what y equals into the second equation 3x + 4 (9 - 2x) = 26 use the distributive property with the 4 3x + 36 - 8x =26 combine like terms -5x + 36 = 26 subtract 36 from both sides -5x = -10 now divide x = 2 now, with x = 2 plug it into the original equation 3 * 2 + 4y = 26 solve like a regular algebra problem 6 + 4y = 26 4y = 20 y = 5

OpenStudy (anonymous):

dang you beat me to it i really need to work on my typing speed

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

in the first sentence i meant you can use elimination or substitution